OpenStudy (anonymous):

4x+5=6-5y solve for x

OpenStudy (anonymous):

inverse operations. do you know how to do this.

OpenStudy (anonymous):

no :O(

OpenStudy (anonymous):

alright. think of the equation as balanced. if you do one thing to one side you have to do it to another. so thinking about inverse operations + - * / derivatives and integrals log and exp ^2 and square roots. and so on. these are inverse operations. you use these when solving equations. now let me know how you think you should solve this.

OpenStudy (anonymous):

4x+5-5=6-5-5y 4x=1-5y 4x/4=1/4&-5y/4 x=1/4-5y/4???

OpenStudy (anonymous):

that's correct

OpenStudy (anonymous):

Thank you so much!!!